# Text-Trained LLMs Can Zero-Shot Extrapolate PDE Dynamics

### v1

@@ 3 Methodology
We extend the tokenization framework introduced by Gruver et al. (2023), which serializes one-dimensional time series as comma-separated numeric strings (e.g., $153,412,\ldots,807$”) for use with LLMs. Our approach, illustrated in Figure 1, generalizes this to time-dependent PDEs by converting discretized spatiotemporal solutions into structured 1D sequences. To encode both spatial and temporal structure, we introduce a two-delimiter format that separates spatial points and time steps. This representation preserves the underlying dynamics and enables interpretable analysis of how LLMs extrapolate higher-dimensional behavior.
@@
Grid Sampling. We begin with a PDE solution $u(x,t)$ evaluated on a regular Cartesian spatiotemporal grid $\{u(x_{i},t_{j})\}_{i=1,j=0}^{N_{\mathrm{X}},N_{\mathrm{T}}}$, represented as a matrix ${\bm{U}}\in\mathbb{R}^{N_{\mathrm{X}}\times(N_{\mathrm{T}}+1)}$. Each column ${\bm{U}}_{:,j}$ corresponds to the spatial state at time $t_{j}$.
@@
Quantization. We apply linear quantization to map the continuous range $u_{\min},u_{\max}]$ to a fixed integer set111000000149149 and 851851999999 are reserved; outputs in these ranges flag explicit out-of-distribution events. $\mathcal{Z}=\{150,151,\dots,850\}\subseteq{\mathbb{Z}}$, yielding a quantized matrix ${\bm{Q}}\in\mathcal{Z}^{N_{\mathrm{X}}\times(N_{\mathrm{T}}+1)}$. This step introduces an underlying quantization error, computed as the difference between the original floating-point values $u(x_{i},t_{j})$ and their concomitant linearly reconstructed counterparts $\tilde{u}(x_{i},t_{j})$, obtained from quantized tokens (see Appendix A.1 for reconstruction details). We refer to this as the quantization floor in subsequent experiments. [^2] Reducible by enlarging $\mathcal{Z}$ beyond its 3-digit representation, at the cost of more tokens per value.
@@
Serialization. Each time slice ${\bm{Q}}_{:,j}$ is serialized into a comma-separated string of 3-digit integers. Temporal evolution is encoded as a sequence of these strings, delimited by semicolons:
@@
We adopt commas to delimit spatial entries, extending the CSV-style format of Gruver et al. (2023) to our setting. To represent the additional temporal dimension in a 2D spatiotemporal matrix, we introduce semicolons to mark time-step boundaries. This enhances parsability and aligns with familiar conventions: semicolons denote row breaks in MATLAB arrays and signal the end of statements in many programming languages (e.g., C, C++, Java). Linguistically, the semicolon also marks a stronger pause than a comma—reinforcing its role as a clear separator between time steps.
@@
Tokenizer Compatibility. We adopt tokenizer configurations (e.g., GPT-4 (Achiam et al., 2023), Llama-3) in which each 3-digit value (000–999) and each delimiter (, and ;) maps to a single token. This one-to-one mapping directly aligns token positions with grid values in discretized PDE solutions, enabling efficient error computation and, importantly, direct estimation of predictive uncertainty at each location from the model’s softmax outputs. In contrast, some models, such as Gemma 3 (Team et al., 2025), tokenize numeric values at different granularities (e.g., each digit as one token). Spatial value probabilities can still be recovered using hierarchical softmax methods in such models (Gruver et al., 2023; Liu et al., 2024), but at a higher computational cost. We therefore focus on LLMs with 3-digit tokenizers, while our serialization remains compatible with other tokenization schemes.
@@
LLM Inference. The serialized sequence is passed to LLMs without fine-tuning or any natural language prompting. Tokens are generated autoregressively using the default generation configuration, with each prediction conditioned on the preceding context. We consider two inference modes: one-step prediction, where the model is set to generate a single future time slice consisting of $2N_{\mathrm{X}}{-}1$ tokens ($N_{\mathrm{X}}$ value tokens and $N_{\mathrm{X}}{-}1$ separator tokens); and multi-step rollouts, which repeat one-step prediction recursively, appending a semicolon after each time slice to indicate temporal progression. At each token position, we record both the model’s output token (for prediction) and its full softmax distribution (for uncertainty analysis). The generated sequence is parsed by splitting on semicolons to segment time and commas to recover spatial locations. [^3] Models rapidly internalize the delimiter structure. Even with minimal context (e.g., one time slice with five spatial points), comma delimiters are consistently generated. Malformed outputs are exceedingly rare, and parsing remains robust across multi-step rollouts. See Section 4.3 for quantitative results.
4 Experiments and Analysis
This section investigates the ability of state-of-the-art open-weight foundation LLMs to continue the spatiotemporal dynamics of PDEs. We focus our main analysis on the Allen–Cahn equation (Allen & Cahn, 1979), a nonlinear PDE modeling phase separation in multi-component metal alloy systems. To assess generality, we additionally evaluate the Fisher–KPP equation (Fisher, 1937; Kolmogorov et al., 1937), another nonlinear PDE modeling population growth and diffusion, together with two representative linear PDEs: the heat equation (a parabolic diffusion model) and the wave equation (a hyperbolic wave propagation model) (Evans, 2010). Across these families, with random initial conditions and varying boundary condition types, we observe qualitatively consistent ICL behavior. Remarkably, LLM rollouts also approximately conserve total thermal energy in the heat equation under Neumann boundaries, indicating that zero-shot ICL captures not only spatiotemporal dynamics but also structural invariants of PDEs. Full results and discussion of these additional PDE experiments are provided in Appendix A.3.
PDE Setup. By coupling reaction and diffusion dynamics, the Allen–Cahn equation induces strong interactions across space and time, making it a challenging yet physically interpretable testbed for assessing whether LLMs capture genuinely spatiotemporal structure rather than merely extrapolating along a single dimension. The system is defined on the interval [1,1][-1,1] with Dirichlet boundary conditions u(1,t)=u(1,t)=1u(-1,t)=u(1,t)=-1 and random initial conditions $u(x,0)=u_{0}(x)$ (details of initial condition generation in Appendix A.1). Explicitly, the Allen–Cahn PDE is:
Here, $\partial_{t}$ and $\partial_{xx}$ denote the temporal and second-order spatial derivatives, respectively. Adopting standard parameter choices (Raissi et al., 2019; Tang & Yang, 2016), we set the diffusion coefficient to $\epsilon^{2}=0.001$ and use a double-well potential $f(u)=2(u^{3}-u)$ for the nonlinear reaction term. The solution is evaluated on a uniform spatiotemporal grid $\smash{\{u(x_{i},t_{j})\}_{i=1,j=0}^{N_{\mathrm{X}},N_{\mathrm{T}}}}$, where $\smash{\{x_{i}\}_{i=1}^{N_{\mathrm{X}}}}$ are $N_{\mathrm{X}}$ evenly spaced interior points and $\smash{\{t_{j}\}_{j=0}^{N_{\mathrm{T}}}}$ are $N_{\mathrm{T}}+1$ evenly spaced time levels, with $T=0.5$.
Numerical Benchmarks. To contextualize the predictive structure learned by LLMs, we compare their outputs to two classical finite difference methods: the forward time, centered space (FTCS) scheme, which is fully explicit, and an implicit-explicit (IMEX) scheme that treats diffusion implicitly and the nonlinear reaction term explicitly (Smith, 1985). In contrast to developing new PDE solvers, our focus is on analyzing how pretrained foundation LLMs extrapolate PDE-based spatiotemporal dynamics in-context, without fine-tuning. Notably, their predictions can achieve surprising accuracy relative to standard numerical benchmarks, generalizing across varied initial conditions and discretizations. These results position PDEs as effective vehicles for examining the inductive biases and generalization behaviors of LLMs.
Analysis Task Overview. We analyze how LLMs generalize in autoregressively continuing PDE dynamics through two analytical lenses: (i) a truncation-error perspective, motivated by local and global error analysis in numerical PDEs (LeVeque, 2007; Larsson & Thomeé, 2003), examining how prediction accuracy depends on discretization, rollout horizon, and model size; and (ii) a “systems-level” perspective, investigating how LLMs internalize and extrapolate PDE structure during ICL via entropy-based uncertainty measures. Section 4.1 analyzes one-step prediction, where accuracy improves with longer temporal context but degrades with finer spatial discretization. Section 4.2 analyzes multi-step rollouts, showing algebraic error growth with rollout horizon, analogous to global error accumulation in numerical solvers. Section 4.3 analyzes token-level uncertainty, revealing a consistent three-stage ICL progression: syntax mimicry, high-entropy exploration, and confident prediction. Prediction quality is evaluated using the Root Mean Square Error (RMSE), computed by aggregating errors over all spatial grid points $\{x_{i}\}_{i=1}^{N_{\mathrm{X}}}$ per time step $\{t_{j}\}_{j=1}^{N_{\mathrm{T}}}$:
RMSE measures overall predictive accuracy. Results using the Maximum Absolute Error, capturing worst-case deviation, show qualitatively similar trends (Appendix A.2). To ensure that our error metrics capture physically meaningful discrepancies—rather than differences between raw 3-digit quantized integer tokens—we compare each predicted solution $\hat{u}(x_{i},t_{j})$ against a floating-point reference solution $\tilde{u}(x_{i},t_{j})$. The reference $\tilde{u}(x_{i},t_{j})$ is computed on a highly refined finite-difference grid and passed through the same quantization–reconstruction pipeline (see Section 3) to ensure a consistent evaluation basis. LLM predictions, generated as token sequences, are likewise reconstructed into floating-point form before error evaluation. Since classical solvers operate on floating-point data, we apply the same quantization–reconstruction process to their initial conditions, ensuring that both LLM-based and classical methods operate on inputs of matched precision. Errors are then averaged over multiple random initial conditions, with further details in Appendix A.2.
Model Setup. Our primary results focus on base pretrained models from the Meta Llama-3 family: Llama-3.1-8B, Llama-3.2-3B, and Llama-3.2-1B. We also evaluate their instruction-tuned counterparts optimized for dialogue in Appendix A.4, which display qualitatively similar trends. For broader comparison, we include Microsoft Phi-4 and Hugging Face SmolLM3, which exhibit similar behaviors with differences mainly in error magnitude, in Appendix A.5.
Figure 2: In-context error scaling with temporal discretization (left) and spatial discretization (right). The top axes show $N_{\mathrm{T}}$ and $N_{\mathrm{X}}$, while the bottom axes show the equivalent LLM context and output lengths $N_{\mathrm{Tokens}}$, respectively. RMSE decreases with longer context, converging in the extended-context regime, toward the local truncation behavior of first-order-in-time solvers (FTCS, IMEX). In contrast, errors grow with output length, following a capacity-dependent generalization trend. Shaded regions show 95% confidence intervals over 50 random initial conditions. The gray dotted line indicates the unavoidable quantization error floor defined in Section 3.
4.1 one-step prediction
We evaluate one-step prediction error under the following setup: given a discretized PDE solution from the initial condition up to one step before the final time, $\{u(x_{i},t_{j})\}_{i=1,j=0}^{N_{\mathrm{X}},N_{\mathrm{T}}-1}$, the model predicts the terminal step $\{u(x_{i},t_{N_{\mathrm{T}}})\}_{i=1}^{N_{\mathrm{X}}}$ across all spatial grid points. This prediction target is employed in order to isolate the effects of temporal and spatial discretization on accuracy. To ensure the task remains non-trivial, we select discretizations where solutions vary significantly between time steps, avoiding degenerate cases in which the model could succeed through trivial pattern repetition. Appendix A.7 provides empirical validation of this design choice.
Effect of Context Length on In-Context Prediction Accuracy. We analyze how one-step prediction accuracy varies with input context length, measured by the number of observed time steps $N_{\mathrm{T}}$ provided to the LLM during ICL. Fixing the spatial discretization at $N_{\mathrm{X}}=14$, we vary $N_{\mathrm{T}}$ from 2 to 40, spanning minimal to moderately informative temporal contexts. As $N_{\mathrm{T}}$ increases, the LLM receives a longer input sequence, and RMSE decreases consistently (Figure 2, left). On a log–log scale, the error curves exhibit decay trends that closely resemble the convergence of first-order-in-time solvers such as FTCS and IMEX as input context length increases, suggesting that LLMs exhibit inductive biases analogous to local truncation error behavior.
Closer examination of Figure 2 (left) reveals three distinct stages as the temporal context increases. The emergence and evolution of these stages—and their connection to prediction uncertainty—are further analyzed in Section 4.3. In the limited-context stage ($N_{\mathrm{T}}=2$), LLMs exhibit substantially higher errors compared to classical solvers. This behavior arises from surface-level pattern imitation in the solution format, rather than learning the underlying dynamical structure (see Section 4.3). In the moderate-context stage ($2<N_{\mathrm{T}}<10$), errors decay more rapidly than those of standard numerical benchmarks such as FTCS and IMEX, suggesting that LLMs move beyond pattern imitation and begin to internalize aspects of the governing PDE dynamics. Finally, in the extended-context stage ($N_{\mathrm{T}}\geq 10$), error decay closely matches that of classical first-order solvers, indicating that LLMs are effectively leveraging the spatiotemporal structure in a numerically grounded way. In this stage, Llama-3.1-8B consistently matches, or in some cases exceeds, the accuracy of classical solvers. Overall, this reveals an empirical in-context scaling law: increasing the input context length consistently improves prediction accuracy, reflecting the LLM’s increasing ability to internalize and extrapolate latent PDE dynamics at fixed spatial discretization.
Effect of Output Length on In-Context Prediction Accuracy. We analyze how prediction accuracy varies with the number of spatial discretization points $N_{\mathrm{X}}$. Since LLMs predict the solution at all spatial grid points for a given time step, increasing $N_{\mathrm{X}}$ directly results in a proportionally longer output sequence. To isolate this effect, we vary $N_{\mathrm{X}}$ across 20 evenly spaced values from 2 to 40, while fixing the temporal context at $N_{\mathrm{T}}=50$. As shown in Figure 2 (right), RMSE grows consistently with $N_{\mathrm{X}}$, following approximate power-law trends on a log–log scale. The error growth is steepest for the smaller Llama-3.2-1B model, while the larger Llama-3.1-8B shows slower growth, indicating improved robustness to output length within the Llama-3 family. This behavior stands in sharp contrast to classical finite-difference solvers, where increasing spatial resolution typically does not raise error under a stable scheme. For LLMs, however, outputs are generated as flat autoregressive token sequences: larger $N_{\mathrm{X}}$ leads to longer, more complex outputs that must be generated without access to the underlying PDE, placing growing demands on the model’s ICL capacity.
These findings reveal a second empirical in-context scaling law: finer spatial discretization produces longer output sequences, which in turn degrade prediction accuracy under fixed input context. The effect scales strongly with model size within the Llama-3 family—smaller models face more pronounced performance drops, indicating that limited ICL capacity constrains generalization at finer spatial discretizations.
4.2 multi-step rollouts
We examine LLMs’ capacity to continue PDE solutions over multiple time steps based solely on in-context input. For a rollout of $N_{\mathrm{T}}$ time steps, we partition the serialized sequence into a context segment of $\lfloor\frac{2}{3}N_{\mathrm{T}}\rfloor$ steps and a prediction segment with the remainder. For $N_{\mathrm{T}}=25$, this yields 16 context steps (including the initial condition), $\{u(x_{i},t_{j})\}_{i=1,j=0}^{N_{\mathrm{X}},15}$, which the LLM uses to autoregressively generate $10$ prediction steps, $\{u(x_{i},t_{j})\}_{i=1,j=16}^{N_{\mathrm{X}},25}$, without access to intermediate ground truth. This 2:1 context-to-prediction ratio strikes a balance between providing sufficient context and posing a nontrivial extrapolation challenge. We assess model behavior via (i) a representative rollout from a single random initial condition, and (ii) average error trends over random initial conditions to quantify how prediction error accumulates over the prediction horizon.
Qualitative Prediction Example from a Single Random Initial Condition.
Figure 3: Multi-step prediction for a randomly sampled initial condition. The left of the dashed line shows the input context provided to the LLM; the right shows a 10-step autoregressive continuation averaged over 20 LLM generations. Classical finite difference solvers (FTCS, IMEX) solve the corresponding initial value problem using the final in-context time slice as the initial condition, and advance the solution for 10 steps using the same spatial and temporal discretization as the LLMs. The second row shows pointwise absolute errors relative to the reference solution.
Figure 3 illustrates a representative multi-step rollout, showing that LLMs can sustain coherent, qualitatively accurate predictions over a 10-step horizon. This is notable because the foundation models are not specialized PDE solvers and lack training-time exposure to the discretized PDE solutions; initial conditions are randomly sampled at inference time. Additional visualizations for other random initial conditions, which exhibit similar behavior, are provided in Appendix A.8. The Llama-3.1-8B model closely tracks the evolution, capturing nonlinear reaction–diffusion dynamics without collapsing to trivial behavior or diverging. When prompted with sufficiently long input context, the model can approximate and extrapolate PDE dynamics purely through autoregressive token-level inference. In contrast, Llama-3.2-1B shows larger deviations and fails to reproduce nonlinear structure over extended rollouts (see Appendix A.9 for analysis of error patterns and capacity limits).
Quantitative Error Trends Across Multiple Random Initial Conditions. To characterize error growth over the prediction horizon, we repeat the multi-step rollout procedure across 20 randomly sampled initial conditions. Averaging over initializations reduces case-specific variability and reveals a consistent trend: RMSE increases algebraically with rollout length, as shown on a log–log scale in Figure 4. This resembles global error accumulation in classical finite-difference solvers operating under stable discretizations, where local truncation errors compound in a controlled manner over time. Crucially, none of the models exhibit divergent or unstable behavior across the 10-step horizon; error growth remains bounded. These findings underscore the capacity of LLMs to continue PDE dynamics via ICL across diverse initial conditions, sustaining coherent predictions over extended horizons—a fundamentally nontrivial task given only in-context information, without prompting or access to governing equations.
Figure 4: Multi-step rollout error trends. RMSE grows algebraically with prediction steps $n$ (top axis) and equivalent LLM output length $N_{\mathrm{Tokens}}$ (bottom axis). Left: rollout from a single random initial condition (as in Figure 3). Right: average over 20 random initial conditions. Error growth rates are estimated via log–log fits and reported on the right. Shaded regions denote 95% confidence intervals (left: across 20 repeated LLM runs; right: across 20 initial conditions).
4.3 Uncertainty Evolution and Learning Stages in In-Context Learning
We now move beyond truncation-like error analysis to examine how LLMs internalize PDE dynamics through ICL. We focus on predictive uncertainty and generation behavior in the one-step prediction task introduced in Section 4.1.444Since the multi-step rollouts analyzed in Section 4.2 recursively apply the one-step prediction process, we defer a parallel analysis of uncertainty accumulation in that setting to Appendix A.10. We vary context length via temporal discretization ($N_{\mathrm{T}}$) and output length via spatial discretization ($N_{\mathrm{X}}$), and analyze how these factors shape the model’s token-level uncertainty. Given input $\{u(x_{i},t_{j})\}_{i=1,j=0}^{N_{\mathrm{X}},N_{\mathrm{T}}-1}$, the model predicts $\{u(x_{i},t_{N_{\mathrm{T}}})\}_{i=1}^{N_{\mathrm{X}}}$. To quantify predictive uncertainty, we compute the Shannon entropy (Shannon, 1948) of the model’s softmax distribution at each spatial value token and average across space:
where $\mathcal{V}$ denotes the tokenizer’s vocabulary, and $p(y\mid x_{i},N_{\mathrm{T}})$ is the predicted probability of token $y$ at spatial location $x_{i}$, given $N_{\mathrm{T}}$ prior time steps in the serialized input. See Appendix A.2 for implementation details.
Emergent Learning Stages Revealed by Entropy Evolution. Figure 5A shows the evolution of mean spatial entropy, $\bar{H}$, as a function of $N_{\mathrm{T}}$. A distinct rise-and-fall pattern reveals three emergent stages of ICL: an initial syntax-dominated stage, a transitional exploratory stage, and a final stage of consolidation and refinement:
  1. 1)
    Syntax-Only (limited context, e.g., $N_{\mathrm{T}}=2$; Figure 5C, first row): While mean spatial entropy $\bar{H}$ exhibits some variation across model sizes, prediction error remains consistently high. Separator tokens (e.g., commas) are predicted with near-perfect confidence (see Figure 5D). In contrast, spatial value tokens act as generic placeholders, with little correspondence to underlying PDE dynamics, yielding deterministic yet physically implausible predictions. This indicates that syntax is acquired before any meaningful understanding of the PDE dynamics emerges.
  2. 2)
    Exploratory (moderate context, e.g., $2<N_{\mathrm{T}}<10$; Figure 5C, second row): Entropy reaches its peak across model sizes, indicating increased uncertainty and broader spatial token distributions. Meanwhile, prediction accuracy improves rapidly, and outputs begin to align with true PDE dynamics. This stage marks a transition from merely capturing surface-level syntax to beginning to internalize spatiotemporal dynamics.
  3. 3)
    Consolidation (extended context, e.g., $N_{\mathrm{T}}\geq 10$; Figure 5C, third row): As context length increases further, $\bar{H}$ decreases, reflecting sharper and more confident spatial token distributions. Prediction accuracy continues to improve, though with less profound gains compared to the exploratory stage. The model’s predictions increasingly reflect coherent and physically meaningful PDE dynamics.
These stages reveal a consistent ICL progression: syntax acquisition, exploratory numerical behavior, and ultimately convergence to accurate predictions. This progression suggests that LLMs develop structured internal representations of PDE dynamics purely through in-context exposure, without explicit access to governing equations or language prompting.
Uncertainty Growth with Output Length. While the previous analysis focused on how input context length affects ICL and predictive uncertainty, we now examine how spatial discretization ($N_{\mathrm{X}}$) affects prediction confidence. Fixing the temporal context at $N_{\mathrm{T}}=50$, we vary $N_{\mathrm{X}}$ and compute the mean spatial entropy $\bar{H}$. As shown in Figure 5B, $\bar{H}$ increases steadily with larger $N_{\mathrm{X}}$, reflecting growing uncertainty for longer spatial outputs. This trend mirrors the error scaling in Figure 2, where RMSE increases with $N_{\mathrm{X}}$ under fixed input context. Notably, the smaller Llama-3.2-1B exhibits the steepest entropy growth, while the larger Llama-3.1-8B shows the slowest, indicating greater robustness to output length. These findings reveal a close empirical link between model uncertainty and prediction error: longer output sequences from finer spatial discretization lead to both higher entropy and reduced predictive accuracy. Within the Llama-3 family, larger models consistently maintain higher confidence and accuracy over extended output lengths. This suggests that model size plays a critical role in enabling accurate extrapolation of learned PDE dynamics across finer spatial discretizations—a capacity that smaller models fail to maintain.
Figure 5: (A–B) Mean spatial entropy $\bar{H}$ vs. (A) temporal context length $N_{\mathrm{T}}$ at fixed $N_{\mathrm{X}}=14$ and (B) output length $N_{\mathrm{X}}$ at fixed $N_{\mathrm{T}}=50$. Shaded regions: 95% confidence intervals over 50 random initial conditions. (C) Token-level softmax distributions at three ICL stages: $N_{\mathrm{T}}=2$ (syntax-only), 55 (exploratory), and 2020 (consolidation), extracted from Llama-3.1-8B for the same initial condition as the multi-step rollout example. Top 8 tokens (by probability) are shown per spatial position; only odd positions are displayed, with full results in Appendix A.11. (D) Softmax over separator tokens, invariant across stages.
5 Conclusion
We show that text-trained LLMs can continue PDE dynamics in-context in a zero-shot setting, without any fine-tuning or natural language prompting. Their performance exhibits clear in-context scaling laws: accuracy improves with longer temporal context, degrades with finer spatial discretization, and error grows algebraically under multi-step rollouts. Entropy analysis further reveals a three-phase progression—from syntax imitation, to exploratory uncertainty, to stabilized predictions—highlighting emergent mechanisms underlying ICL. Together, these findings suggest that LLMs can internalize nontrivial aspects of PDE dynamics purely from in-context data, demonstrating emergent generalization capabilities in zero-shot inference.
Limitations and Future Work. Our study is restricted to time-dependent PDEs with real-valued solutions on regular domains. Extending this framework to stationary PDEs (e.g., Poisson), PDEs with complex-valued solutions (e.g., Schrödinger), higher-dimensional or irregular domains, and partially observed or noisy dynamics could reveal complementary behaviors. Beyond these extensions, a key challenge is to characterize the internal representations and compositional structures that support generalization over spatiotemporal dynamics in autoregressive token space. While we focus on demonstrating the surprising ICL abilities of text-trained LLMs for PDE dynamics continuation, an important open problem is to identify PDE regimes where LLMs outperform classical solvers, and to explore how symbolic priors or textual prompts (Xue & Salim, 2024; Requeima et al., 2024) could amplify these advantages.
Appendix A Appendix
A.1 Details on Experimental Setup and Initial Conditions
A.1.1 Quantization and Reconstruction Implementation

### v2

@@ Methods
We extend the tokenization framework introduced by Gruver et al.[9], which serializes one-dimensional time series as comma-separated numeric strings (e.g., $153,412,\ldots,807$”) for use with LLMs. Our approach, illustrated in Fig. 1, generalizes this to time-dependent PDEs by converting discretized spatiotemporal solutions into structured 1D sequences. To encode both spatial and temporal structure, we introduce a two-delimiter format that separates spatial points and time steps. This representation preserves the underlying dynamics and enables interpretable analysis of how LLMs extrapolate higher-dimensional behavior.
Grid Sampling.
@@
We begin with a PDE solution $u(x,t)$ evaluated on a uniform spatiotemporal grid $\{u(x_{i},t_{j})\}_{i=1,j=0}^{N_{\mathrm{X}},N_{\mathrm{T}}}$, represented as a matrix $\bm{U}\in\mathbb{R}^{N_{\mathrm{X}}\times(N_{\mathrm{T}}+1)}$. Each column $\bm{U}_{:,j}$ corresponds to the spatial state at time $t_{j}$. evaluations on non-uniform spatiotemporal grids are provided in Supplementary Material Sect. 8, which exhibit qualitatively similar behavior to the uniform-grid results.
Quantization.
@@
We apply linear quantization to map the continuous range $u_{\min},u_{\max}]$ to a fixed integer set $\mathcal{Z}=\{150,151,\dots,\allowbreak 850\}\subseteq\mathbb{Z}$, yielding a quantized matrix $\bm{Q}\in\mathcal{Z}^{N_{\mathrm{X}}\times(N_{\mathrm{T}}+1)}$. The ranges 000000149149 and 851851999999 are reserved to flag explicit out-of-distribution events. This step introduces an underlying quantization error, computed as the difference between the original floating-point values $u(x_{i},t_{j})$ and their concomitant linearly reconstructed counterparts $\tilde{u}(x_{i},t_{j})$, obtained from quantized tokens (see Supplementary Material Sect. 1 for reconstruction details). We refer to this as the quantization floor in subsequent experiments, which can be reduced by enlarging $\mathcal{Z}$ beyond a 3-digit representation, at the cost of more tokens per value.
Serialization.
@@
Each time slice $\bm{Q}_{:,j}$ is serialized into a comma-separated string of 3-digit integers. Temporal evolution is encoded as a sequence of these strings, delimited by semicolons:
@@
We adopt commas to delimit spatial entries, extending the CSV-style format of Gruver et al[9] to our setting. To represent the additional temporal dimension in a 2D spatiotemporal matrix, we introduce semicolons to mark time-step boundaries. This enhances parsability and aligns with familiar conventions: semicolons denote row breaks in MATLAB arrays and signal the end of statements in many programming languages (e.g., C, C++, Java). Linguistically, the semicolon also marks a stronger pause than a comma, reinforcing its role as a clear separator between time steps.
Tokenizer Compatibility.
@@
We adopt tokenizer configurations (e.g., GPT-4 [40], Llama-3) in which each 3-digit value (000–999) and each delimiter (, and ;) maps to a single token. This one-to-one mapping directly aligns token positions with grid values in discretized PDE solutions, enabling efficient error computation and, importantly, direct estimation of predictive uncertainty at each location from the model’s softmax outputs. In contrast, some models, such as Gemma 3 [41], tokenize numeric values at different granularities (e.g., each digit as one token). Spatial value probabilities can still be recovered using hierarchical softmax methods in such models [9, 11], but at a higher computational cost. We therefore focus on LLMs with 3-digit tokenizers, while our serialization remains compatible with other tokenization schemes.
LLM Inference.
@@
The serialized sequence is passed to LLMs without fine-tuning or any natural language prompting. Tokens are generated autoregressively using the default generation configuration, with each prediction conditioned on the preceding context. We consider two inference modes: one-step prediction, where the model is set to generate a single future time slice consisting of $2N_{\mathrm{X}}{-}1$ tokens ($N_{\mathrm{X}}$ value tokens and $N_{\mathrm{X}}{-}1$ separator tokens); and multi-step rollouts, which repeat one-step prediction recursively, appending a semicolon after each time slice to indicate temporal progression. At each token position, we record both the model’s output token (for prediction) and its full softmax distribution (for uncertainty analysis). The generated sequence is parsed by splitting on semicolons to segment time and commas to recover spatial locations. Models rapidly internalize the delimiter structure. Even with minimal context (e.g., one time slice with five spatial points), comma delimiters are consistently generated. Malformed outputs are exceedingly rare, and parsing remains robust across multi-step rollouts. This behavior is quantified in the Uncertainty Evolution and Learning Stages section.
Data availability
All data and Python code to reproduce the numerical experiments in the main text and the Supplementary Material are publicly available in the GitHub repository at https://github.com/Jiajun-Bao/LLM-PDE-Dynamics.
Supplementary Material for
Text-Trained LLMs Can Zero-Shot Extrapolate PDE Dynamics, Revealing a Three-Stage In-Context Learning Mechanism
Contents
1 Details on Experimental Setup and Initial Conditions
1.1 Quantization and Reconstruction Implementation